# Precise quantum-geometric electronic properties from first principles

### v1

@@ Abstract
The calculation of quantum-geometric properties of Bloch electrons – Berry curvature, quantum metric, orbital magnetic moment and effective mass – was implemented in a pseudopotential plane-wave code. The starting point was the first derivative of the periodic part of the wavefunction $\psi({\bf k})$ with respect to the wavevector ${\bf k}$. This was evaluated with perturbation theory by solving a Sternheimer equation, with special care taken to deal with degenerate levels. Comparison of effective masses obtained from perturbation theory for silicon and gallium arsenide with carefully-converged numerical second derivatives of band energies confirms the high precision of the method. Calculations of quantum-geometric quantities for gapped graphene were performed by adding a bespoke symmetry-breaking potential to first-principles graphene. As the two bands near the opened gap are reasonably isolated, the results could be compared with those obtained from an analytical two-band model, allowing to assess the strengths and limitations of such widely-used models. The final application was trigonal tellurium, where quantum-geometric quantities flip sign with chirality.
@@ 1 Introduction
Exploring the quantum-geometric and topological properties of electrons in crystals is a very active research topic [1, 2, 3]. Band effective masses, which have been a staple of semiconductor physics since the early days of solid-state physics, are now recognized as one of the basic quantum-geometric properties of materials, along with the Berry curvature, quantum metric, and intrinsic orbital moment. A characteristic of such quantities is that they can vary by several orders of magnitude across the Brillouin zone (BZ), with large absolute values usually associated with small regions near band degeneracies or quasi-degeneracies. Therefore, a very fine sampling of the BZ may be required to find those regions, or to calculate properties that are given by BZ integrals of expressions involving quantum-geometric quantities [4]. Interpolation methods, in particular Wannier interpolation, have been extensively used to calculate such properties [5, 6, 7], although the evaluation of the quantum metric was implemented only recently [8]. In some cases – for example, the band edges of semiconductors – the region of interest in the BZ may be known, allowing the calculation to be done without recourse to an interpolation scheme.
@@
The paper is organized as follows. In Sec. 2 the context, method and implementation are presented, paying special attention to the treatment of band degeneracies. To illustrate the method, the precise effective masses in centrosymmetric silicon and noncentrosymmetric gallium arsenide are presented in Sec. 3; the convergence of the finite-differences and ${\bf k}\cdot{\bf p}$ methods is also discussed, using our perturbative implementation as a reference. In Sec. 4 deals with an artificial system obtained by breaking the inversion symmetry of graphene, thereby opening a gap at the Dirac points. This allows for a detailed comparison of the calculated ab initio quantum-geometric quantities with analytical results from a two-band effective model. The final example, in Sec. 5, is trigonal tellurium, a chiral semiconducting crystal with striking electronic responses driven by quantum geometry at the band edge [16, 17, 18, 19, 20, 21, 22]. The article concludes in Sec. 6 with a summary.
@@ 2.1 Periodic potential
with one-electron energy levels $E_{n}$ and wavefunctions $\psi_{dn}({\bf r})$. The energies are assumed to be in strictly increasing order, $E_{1}<E_{2}<E_{3}<\ldots$, and each energy level has degeneracy $D_{n}$, with $d=1,2,\ldots,D_{n}$ an aditional wavefunction index. Wavefunctions within a degenerate subspace are chosen to be orthonormal: $\langle\psi_{dn}|\psi_{d^{\prime}m}\rangle=\delta_{dd^{\prime}}\delta_{nm}$. Here and throughout the paper, we use atomic units ($\hbar=m_{e}=|e|=4\pi\epsilon_{0}=1$) to simplify the notation, unless stated otherwise.
@@
The Bloch wavefunction can be expressed as the product of a periodic function $u_{dn{\bf k}}({\bf r})$ and a spatial phase,
@@
The weights $w_{j}$ can be positive or negative, and may be assumed to be just ±$\pm 1$ if the KB projectors are not normalized.
@@ 2.2 Perturbation theory
$({E}_{n}-H_{\bf k})Q_{n{\bf k}}|{u}_{dn{\bf k}}^{\alpha}\rangle=Q_{n{\bf k}}H_{\bf k}^{\alpha}|{u}_{dn{\bf k}}\rangle.$ (13)
Thanks to the disappearance of the highly nonlinear first-order energy term $E_{dn{\bf k}}^{\alpha}$ from the right-hand side, this is now a linear equation; therefore, any linear combination of the right-hand side will have as a solution the same linear combination of wavefunction perturbations. This means that once solutions have been found for three linearly independent directions, the solution for any direction can be recovered.
@@ 2.3 Band velocities
whose eigenvalues $E_{jn{\bf k}}^{(1)}(\hat{\bf q})$ are the band velocities along that particular direction. (Equation (21) is the matrix representation of the first-order Hamiltonian $H_{\bf k}^{\hat{\bf q}}=\sum_{\alpha}\hat{q}_{\alpha}H_{\bf k}^{\alpha}$ within the degenerate subspace.)
@@ 2.4 Quantum-geometric quantities
Once we have $H_{\bf k}$, $H_{\bf k}^{\alpha}$, $H_{\bf k}^{\alpha\beta}$, and $Q_{n{\bf k}}|{u}_{dn{\bf k}}^{\alpha}\rangle$, we can calculate several quantum-geometric quantities, including effective masses, with just algebraic operations. Given a group of $D_{n\bf k}$ degenerate bands at ${\bf k}$, one can associate with it a quantum-geometric tensor defined as [2, 3, 13]
@@
while the Cartesian symmetric (or band Hermitian) part contributes to a generalized effective mass tensor of the band group, which can be conveniently written as follows,
@@
(Here we restored the constants $\hbar$ and $|e|$ for clarity.) For a nondegenerate level, the orbital magnetic moment and effective mass tensors become [33, 1, 34]
@@ 2.5 Calculation of effective masses in a given direction
$U^{\dagger}A^{(1)}U=E_{n{\bf k}}^{(1)}.$ (34)
Here, $E_{n{\bf k}}^{(1)}$ is a diagonal matrix with ordered diagonal elements $E_{1n{\bf k}}^{(1)}(\hat{\bf q})<E_{2n{\bf k}}^{(1)}(\hat{\bf q})<\ldots<E_{J_{n{\bf k}}n{\bf k}}^{(1)}(\hat{\bf q})$. There are $J_{n{\bf k}}(\hat{\bf q})$ distinct eigenvalues of $A^{(1)}(\hat{\bf q},n,{\bf k})$, each with degeneracy $I_{jn{\bf k}}(\hat{\bf q})$, satisfying $\sum_{j=1}^{J_{n{\bf k}}(\hat{\bf q})}I_{jn{\bf k}}(\hat{\bf q})=D_{n{\bf k}}$. The same transformation must now be applied to the second-order matrix $A_{dd^{\prime}}^{(2)}(\hat{\bf q},n,{\bf k})$,
@@
yielding a new second-order matrix ${\tilde{A}}_{dd^{\prime}}^{(2)}(\hat{\bf q},n,{\bf k})$. Finally, for each band velocity $E_{jn{\bf k}}^{(1)}$ we extract from the associated rows and columns of ${\tilde{A}}_{dd^{\prime}}^{(2)}(\hat{\bf q},n,{\bf k})$ the submatrix ${\tilde{\tilde{A}}}_{dd^{\prime}}^{(2)}(\hat{\bf q},j,n,{\bf k})$ of dimensions $I_{jn{\bf k}}\times I_{jn{\bf k}}$. The eigenvalues $E_{ijn{\bf k}}^{(2)}(\hat{\bf q})$ of ${\tilde{\tilde{A}}}_{dd^{\prime}}^{(2)}(\hat{\bf q},j,n,{\bf k})$ are the inverse effective masses along $\hat{\bf q}$,
$\left.\frac{\partial^{2}}{\partial\eta^{2}}E_{n,{\bf k}+\eta\hat{\bf q}}(i,j)\right|_{\eta=0}=E_{ijn{\bf k}}^{(2)}(\hat{\bf q}).$ (36)
We note that the procedure outlined above is equivalent to applying the unitary transformation $U$ that diagonalizes $A^{(1)}$ to the original wavefunctions $|{u}_{jn{\bf k}}(\hat{\bf q})\rangle$ to obtain a new set of wavefunctions $|{\tilde{u}}_{ijn{\bf k}}(\hat{\bf q})\rangle$, with $i=1,\ldots,I_{jn{\bf k}}(\hat{\bf q})$ and $j=1,\ldots,J_{n\bf k}(\hat{\bf q})$; the new set is then used to recalculate Eqs. (28) and (30), as they are the relevant wavefunctions for carrying out second-order degenerate perturbation theory [27] in the subspace with first-order eigenvalue corrections $E_{jn{\bf k}}^{(1)}(\hat{\bf q},n,{\bf k})$. However, applying the transformations directly to the operators is computationally more efficient.
Effective masses of degenerate bands have been calculated from first principles using perturbative methods (as opposed to finite differences) with the abinit [9] and WIEN2k [12] codes; for the nondegenerate case, there is earlier work [36]. As those calculations, as well as the present work, use perturbation theory to calculate effective masses, and their basis sets are all related to plane waves – linearized augmented-plane-wave (LAPW), projected augmented wave (PAW) or just elementary plane waves – there are close similarities between all three implementations. The main distinction between the present approach and the previous work is that we use $Q_{n{\bf k}}|{u}_{dn{\bf k}}^{\beta}\rangle$ given by the Sternheimer equation (13) as our basic ingredient, whereas the previous work is based on explicit sums over states; the connection between the two approaches is Eq. (14). The abinit implementation evaluates the effective masses in a plane-wave basis via the expression in Eq. (66) of Ref. [9]. That expression is different from our Eq. (30) with $\Gamma_{dd^{\prime}}^{\alpha\beta}(n,{\bf k})$ defined by Eq. (28), although it is possible to show they are equivalent using Eq. (14). As the abinit implementation uses a Sternheimer equation to compute the sum over states, there are also points in common at the algorithmic level. The present approach has the advantage of yielding the Berry curvature, quantum metric and orbital moment for an almost negligible extra computational cost, presenting the calculation of quantum-geometric quantities in a unified framework.
@@ 3.1 Silicon
The self-consistent potential of Si was calculated in the local-density approximation (LDA) with the Perdew-Wang parametrization [37], using the cpw2000 pseudopotential plane-wave code [15]. The pseudopotential was relativistic Troullier-Martins [38, 39] with a core radius of 1.8 atomic units, ground-state configuration, and $s$, $p$, and $d$ channels. The local potential was a smoothed maximum of all channels, and the nonlocal part was converted to the KB form. The self-consistent calculation used a kinetic energy cutoff of 20 Ha for the plane-wave expansion, and a $6\times 6\times 6$ uniform grid for BZ integration (28 ${\bf k}$ points once symmetry was taken into account). The lattice constant in the calculations was $5.4015\,\,\text{\AA}$. The lattice constant, energy cutoff, exchange-correlation functional, and BZ sampling were chosen to be the same as those of Ref. [9], so that any differences in results could be attributed to the difference in electronic-structure method: projected augmented wave (PAW) in Ref. [9], versus pseudopotential in the present work. The calculations were well converged: increasing cutoff or BZ sampling had typically a minor effect on the 4th decimal place of the effective masses. Using other LDA parametrizations, or using a different core radius, also only had an effect on the 4th decimal place. Using a generalized gradient approximation (GGA) only slightly changed the third decimal place. However, using the experimental lattice constant changed the effective masses by a few percent. As we will be checking numerical accuracy, in some tables we report a high number of decimal places.
@@
Figure 1: The dependence on direction of the second derivative of the band energy (inverse effective mass) is shown for the top of the valence band of Si. The distance from the center of each figure to the surface is proportional to the second derivative of the band energy in that direction. The inverse effective masses are clearly anisotropic, and cubic symmetry is respected. The scale of the two figures is not the same.
Table 1 shows the effective band mass in Si at the $\Gamma$ point for high symmetry directions. States usually described as holes appear with negative masses. States are identified by symmetry, with the added subscript $v$ for valence or $c$ for conduction. The first column contains the results of the present work. They agree within a few percent with the PAW calculations shown in the second column. Effective masses can also be calculated by finite differences from energy bands calculated along a few points in the chosen direction, and we show in that table the results from Lagrange interpolations of order $n=6$ and $8$, with $n+1$ points centered at $\Gamma$ and spaced by $\delta=10^{-3}$, $10^{-4}$ or $10^{-5}$ times $2\pi/a$. Instead of an analytical finite-difference formula for the second derivatives, we use a recursive algorithm, that has the advantage of being numerically stable. As an added bonus, by comparing the second derivatives at order $n-1$ of the recursion with the final value, an indication of the reliability of the estimate is obtained [40]. One can see that for $\delta=10^{-4}$, the $n=6$ and $n=8$ interpolations give the same values for the effective masses as perturbation theory. For the larger $\delta=10^{-3}$ spacing of the sampling wavevectors, there are some deviations in mass values, but they are still accurate for the physics of the problem. For the smaller $\delta=10^{-5}$ spacing, the numerical instability of Lagrange interpolations for closely-spaced points appears, and the results are less reliable. In fact, for that small spacing, the degeneracy of the mass values (they should all be doubly degenerate) is numerically broken, and the values reported in the table are their mean values. The main point of the table is that the implemented method works and is very stable and accurate.
@@
There is an approximation from the early days of electronic structure theory, the ${\bf k}\cdot{\bf p}$ method, that uses the second-order expansion of the Hamiltonian implicit in Eqs. 1, 8 and 9. (Obviously, the early work did not have the projectors of the non-local pseudopotential.) Given a reference wavevector ${\bf k}_{0}$ and its eigenfunctions $|{u}_{dn{\bf k}_{0}}\rangle$, one can construct an approximate Hamiltonian matrix for a neighboring wave vector ${\bf k}$
@@ 3.2 Gallium arsenide
While the general features of the gallium arsenide band structure are similar to those of silicon, the absence of spatial inversion symmetry means that the bands are not necessarily degenerate, and at the top of the valence band at $\Gamma$, the band slope may be non-zero. These effects are easily missed in a band structure plot showing the full energy range of the valence and lower conduction bands, see Fig 2, or even in the scale of room temperature; however, they are clearly seen on the $10^{-5}\,\,\text{eV}$ energy range of the insets of Fig 2. The perturbation theory at the $\Gamma$ point will give the first and second derivatives of the energy at that point, which is connected with the smallest of the energy scales. This behavior near the top of the valence band of GaAs was described by E. Dresselhaus in a seminal paper [42].
Figure 2: The valence bands (blue) and conduction bands (red) of GaAs are shown for the main symmetry directions of the fcc lattice. The highly magnified behavior of the hole bands near the top of the valence band is shown in the insets, where the energy range is $10^{-5}\,\,\text{eV}$ and the wavevector range is $1.18\times 10^{-3}$ atomic units.
The self-consistent potential of GaAs was calculated with the cpw2000 pseudopotential plane-wave code using a modified Becke-Johnson meta-GGA functional [43] that gives an energy gap close to experiment. Both Ga and As pseudopotentials were relativistic, constructed with the Troullier-Martins recipe [38] from a ground-state configuration and LDA exchange and correlation. The Ga pseudopotential had a core radius of 2.5 atomic units for the $s$ and $p$ channels, and 2.69 atomic units for the $d$ channel. The As pseudopotential had a core radius of 2.1 atomic units for the $s$, $p$ and $d$ channels. In both cases the $d$ channel was used as the local part, and the nonlocal part was converted to the KB form. The self-consistent calculation used a kinetic energy cutoff of 16 Ha for the plane-wave expansion, and a $4\times 4\times 4$ uniform grid for BZ integration (10 ${\bf k}$ points once symmetry was taken into account). The lattice constant in the calculations was $5.653\,\,\text{\AA}$.
@@
The calculated effective masses at $\Gamma$ are shown in Table 3. There is an excellent agreement (about six decimal places) between the masses calculated from perturbation theory and those obtained from an interpolation of band energies (with a sixth order polynomial, and points equally spaced by $10^{-5}$ and centered in $\Gamma$). However, in the interpolation, the bands have to be carefully “disentangled” to follow the correct branch after each crossing. Even with the help of the quality estimate of the fit that flags most mistakes [40], that disentanglement is a very tedious and error prone job.
A clear sign of the peculiarities in the effective masses is that at $\Gamma$ and in the $[100]$ direction they are identical for the “heavy” and “light” holes. In that direction, the four bands that are degenerate at $\Gamma$ split into two doubly degenerate bands that are mirror images of each other, and therefore have the same second derivatives at $\Gamma$. From the insets of Fig. 2, it is clear that near $\Gamma$ the bands deviate from parabolas. In Fig. 3, the longitudinal masses along $[100]$ are plotted as a function of distance from the $\Gamma$ point. The reciprocal vector is on a scale $x=log(1+k/k_{0})$ that is linear near the origin ($k\ll k_{0}$) but logarithmic far from it ($k\gg k_{0}$). The mass of the split-off hole is almost constant throughout the plot range. The heavy and light hole masses are identical at $\Gamma$, deviating linearly at very small $k$ until they settle at an almost constant value.

### v2

@@ Abstract
The calculation of quantum-geometric properties of Bloch electrons – Berry curvature, quantum metric, orbital magnetic moment and effective mass – was implemented in a pseudopotential plane-wave code. The starting point was the first derivative of the periodic part of the wavefunction $\psi_{\bf k}({\bf r})$ with respect to wavevector $\bf k$}. This was evaluated with perturbation theory by solving a Sternheimer equation. Comparison of effective masses obtained from perturbation theory for silicon and gallium arsenide with carefully-converged numerical second derivatives of band energies confirmed the high precision of the method. Calculations of quantum-geometric quantities for gapped graphene were performed by adding a bespoke symmetry-breaking potential to first-principles graphene. As the two bands near the opened gap are reasonably isolated, the results could be compared with those obtained from an analytical two-band model, allowing to assess the strengths and limitations of such widely-used models. The final application was trigonal tellurium, where some quantum-geometric quantities flip sign with chirality.
@@ 1 Introduction
Exploring the quantum-geometric and topological properties of electrons in crystals is a very active research topic [1, 2, 3, 4]. Band effective masses, which have been a staple of semiconductor physics since the early days of solid-state physics, are now recognized as one of the basic quantum-geometric properties of materials, along with the Berry curvature, quantum metric, and intrinsic orbital moment. A characteristic of such quantities is that they can vary by several orders of magnitude across the Brillouin zone (BZ), with large absolute values usually associated with small regions near band degeneracies or quasi-degeneracies. Therefore, a very fine sampling of the BZ may be required to find those regions, or to calculate properties that are given by BZ integrals of expressions involving quantum-geometric quantities [5]. Interpolation methods, in particular Wannier interpolation, have been extensively used to calculate such properties [6, 7, 8], although the evaluation of the quantum metric in this manner was implemented only recently [9]. In some cases – for example, the band edges of semiconductors – the region of interest in the BZ may be known, allowing the calculation to be done without recourse to an interpolation scheme.
@@
The paper is organized as follows. In Sec. 2 the context, method and implementation are presented. Special attention is paid to the treatment of band degeneracies, as degenerate perturbation theory is much more complex to implement, and the physical tensors have degeneracy indices added to the spatial indices. To illustrate the method, the precise effective masses in centrosymmetric silicon and noncentrosymmetric gallium arsenide are presented in Sec. 3; the convergence of the finite-differences and ${\bf k}\cdot{\bf p}$ methods is also discussed, using our perturbative implementation as a reference. Section 4 deals with an artificial system obtained by breaking the inversion symmetry of graphene, thereby opening a gap at the Dirac points. This allows for a detailed comparison of the calculated ab initio quantum-geometric quantities with analytical results from a two-band effective model. The final example, in Sec. 5, is trigonal tellurium, a chiral semiconducting crystal with striking electronic responses driven by quantum geometry at the band edge [17, 18, 19, 20, 21, 22, 23]. The article concludes in Sec. 6 with a summary.
@@ 2.1 Periodic potential
with one-electron energy levels $E_{n}$ and wavefunctions $\psi_{dn}({\bf r})$. The energies are assumed to be in strictly increasing order, $E_{1}<E_{2}<E_{3}<\ldots$, and each energy level has degeneracy $D_{n}$, with $d=1,2,\ldots,D_{n}$ an additional wavefunction index. Wavefunctions within a degenerate subspace are chosen to be orthonormal: $\langle\psi_{dn}|\psi_{d^{\prime}m}\rangle=\delta_{dd^{\prime}}\delta_{nm}$, where the normalization volume is one crystal cell. Here and throughout the paper, we use Hartree atomic units ($\hbar=m_{e}=|e|=4\pi\epsilon_{0}=1$) to simplify the notation, unless stated otherwise.
@@
The Bloch wavefunction can be expressed as the product of a periodic function $u_{dn{\bf k}}({\bf r})$ and a spatial phase factor,
@@
The weights $w_{j}$ can be positive or negative, and may be assumed to be just $\pm 1$ if the KB projectors are not normalized.
@@ 2.2 Perturbation theory
$({E}_{n{\bf k}}-H_{\bf k})Q_{n{\bf k}}|{u}_{dn{\bf k}}^{\alpha}\rangle=Q_{n{\bf k}}H_{\bf k}^{\alpha}|{u}_{dn{\bf k}}\rangle.$ (13)
Thanks to the disappearance of the highly nonlinear first-order energy term $E_{dn{\bf k}}^{\alpha}$ from the right-hand side, this is now a linear equation; therefore, any linear combination of the right-hand side will have as a solution the same linear combination of wavefunction perturbations. This means that once solutions have been found for three linearly-independent directions, the solution for any direction can be recovered.
@@ 2.3 Band velocities
[…]
@@ 2.4 Quantum-geometric quantities
Once we have $H_{\bf k}$, $H_{\bf k}^{\alpha}$, $H_{\bf k}^{\alpha\beta}$, and $Q_{n{\bf k}}|{u}_{dn{\bf k}}^{\alpha}\rangle$, we can calculate several quantum-geometric quantities, including effective masses, with just algebraic operations. Given a group of $D_{n\bf k}$ degenerate states at $\bf k$}, one can associate with it a quantum-geometric tensor defined as [2, 3, 14]
@@
while the Cartesian symmetric (or band Hermitian) part contributes to a generalized inverse effective mass tensor of the band group, which can be conveniently written as follows,
@@
(Here we restored the constants $\hbar$ and $|e|$ for clarity.) For a nondegenerate level, the orbital magnetic moment and inverse effective mass tensors become [34, 1, 35]
@@
Figure 1: A flowchart of the procedure for calculating the quantum-geometric quantities and transport equivalent effective masses. The draft user guide, with the instructions to perform these calculations and interpret the results, is available from the same website as the code [16].
@@ 2.5 Calculation of effective masses in a given direction
$U^{\dagger}A^{(1)}U=E^{(1)}.$ (34)
[…]
@@
[…]
A schematic flowchart of the procedure is shown in Fig. 1.
[…]
Effective masses of degenerate bands have been calculated from first principles using perturbative methods (as opposed to finite differences) with the abinit [10] and WIEN2k [13] codes; for the nondegenerate case, there is earlier work [37]. As those calculations, as well as the present work, use perturbation theory to calculate effective masses, and their basis sets are all related to plane waves – linearized augmented-plane-wave (LAPW), projected augmented wave (PAW) or just elementary plane waves – there are close similarities between all three implementations.
The main distinction between the present approach and previous works is that we use $Q_{n{\bf k}}|{u}_{dn{\bf k}}^{\alpha}\rangle$ given by the Sternheimer equation (13) as our basic ingredient, whereas the previous works start from explicit sums over states; the connection between the two approaches is Eq. (14). The abinit implementation evaluates the effective masses in a plane-wave basis via the expression in Eq. (66) of Ref. [10]; that expression is different from our Eq. (30) with $\Gamma_{dd^{\prime}}^{\alpha\beta}(n,{\bf k})$ defined by Eq. (28), although it is possible to show they are equivalent using Eq. (14). As the abinit implementation uses a Sternheimer equation to circunvent an infinite sum over states, there are also points in common at the algorithmic level. The present approach has the advantage of yielding the Berry curvature, quantum metric and orbital moment for an almost negligible extra computational cost, presenting the calculation of quantum-geometric quantities in a unified framework.
@@ 3.1 Silicon
The self-consistent potential of Si was calculated in the local-density approximation (LDA) with the Perdew-Wang parametrization [38], using the cpw2000 pseudopotential plane-wave code [16]. The pseudopotential was relativistic Troullier-Martins [39, 40] with a core radius of 1.8 atomic units, ground-state configuration, and $s$, $p$ and $d$ channels. The local potential was a smoothed maximum of all channels, and the nonlocal part was converted to the KB form. The self-consistent calculation used a kinetic energy cutoff of 20 Ha for the plane-wave expansion, and a $6\times 6\times 6$ uniform grid for BZ integration (28 $\bf k$} points, once symmetry was taken into account). The lattice constant in the calculations was $5.4015\,\,\text{\AA }$. The lattice constant, energy cutoff, exchange-correlation functional, and BZ sampling were chosen to be the same as those of Ref. [10], so that any differences in results could be attributed to the difference in electronic-structure method: PAW in Ref. [10], versus pseudopotential in the present work. The calculations were well converged: increasing cutoff or BZ sampling had typically a minor effect on the 4th decimal place of the effective masses. Using other LDA parametrizations, or using a different core radius, also only had an effect on the 4th decimal place. Using a generalized gradient approximation (GGA) only slightly changed the third decimal place. However, using the experimental lattice constant changed the effective masses by a few percent. As we will be checking numerical accuracy, in some tables we report a high number of decimal places.
@@
Figure 2: The dependence on direction of the second derivative of the band energy (inverse effective mass) is shown for the top of the valence band of Si at the $\Gamma$ point. The distance from the center of each figure to the surface is proportional to the second derivative of the band energy in that direction. The inverse effective masses are clearly anisotropic, and cubic symmetry is respected. The scale of the two figures is not the same.
Table 1 shows the effective band mass in Si at the $\Gamma$ point for high-symmetry directions. States usually described as holes appear with negative masses. States are identified by symmetry, with the added subscript $v$ for valence or $c$ for conduction. The first column contains the results of the present work. They agree within a few percent with the PAW calculations shown in the second column. Effective masses can also be calculated by finite differences from energy bands calculated along a few points in the chosen direction, and we show in that table the results from Lagrange interpolation of order $n=6$ and $8$, with $n+1$ points centered at $\Gamma$ and spaced by $\delta=10^{-3}$, $10^{-4}$ or $10^{-5}$ (in atomic units). Instead of an analytical finite-difference formula for the second derivatives, we use a recursive algorithm, which has the advantage of being numerically stable. As an added bonus, by comparing the second derivatives at order n1n-1 of the recursion with the final value, an indication of the reliability of the estimate is obtained [41]. One can see that for $\delta=10^{-4}$, the $n=6$ and $n=8$ interpolations give the same values for the effective masses as perturbation theory. For the larger $\delta=10^{-3}$ spacing of the sampling wavevectors, there are some deviations in mass values, but they are still sufficiently accurate for the physics of the problem. For the smaller $\delta=10^{-5}$ spacing, the numerical instability of Lagrange interpolation with closely-spaced points appears, and the results are less reliable. In fact, for that small spacing, the degeneracy of the mass values (they should all be doubly degenerate) is numerically broken, and the values reported in the table are their mean values. The main point of the table is that the implemented perturbative method works and is very stable and accurate.
@@
There is an approximation from the early days of electronic structure theory, the ${\bf k}\cdot{\bf p}$ method, that uses the second-order expansion of the Hamiltonian implicit in Eqs. (1), (8) and (9). (Obviously, the early work did not have the projectors of the nonlocal pseudopotential.) Given a reference wavevector ${\bf k}_{0}$ and its cell-periodic eigenfunctions $|{u}_{dn{\bf k}_{0}}\rangle$, one can construct an approximate Hamiltonian matrix for a neighboring wavevector $\bf k$},
@@ 3.2 Gallium arsenide
While the general features of the gallium arsenide band structure are similar to those of silicon, the absence of spatial inversion symmetry means that the bands are not necessarily degenerate, and at the top of the valence band at $\Gamma$, the band slope may be non-zero. These effects are easily missed in a band-structure plot showing the full energy range of the valence and lower conduction bands, see Fig 3, or even on the scale of room temperature; however, they are clearly seen on the $10^{-5}\,\,\text{eV}$ energy range of the insets in Fig 3. Perturbation theory at the $\Gamma$ point will give the first and second derivatives of the energy at that point, which are connected with the smallest of the energy scales. This behavior near the top of the valence band of GaAs was described by E. Dresselhaus in a seminal paper [43].
Figure 3: The valence bands (blue) and conduction bands (red) of GaAs are shown for the main symmetry directions of the fcc lattice. The highly magnified behavior of the hole bands near the top of the valence band is shown in the insets, where the energy range is $10^{-5}\,\,\text{eV}$ and the wavevector range is $1.18\times 10^{-3}$ atomic units. The magenta lines connect the insets to the regions of the main plot that were magnified.
The self-consistent potential of GaAs was calculated with the cpw2000 pseudopotential plane-wave code using a modified Becke-Johnson meta-GGA functional [44] that gives an energy gap close to experiment. Both Ga and As pseudopotentials were relativistic, constructed with the Troullier-Martins recipe [39, 40] from a ground-state configuration and LDA exchange and correlation. The Ga pseudopotential had a core radius of 2.5 atomic units for the $s$ and $p$ channels, and 2.69 atomic units for the $d$ channel. The As pseudopotential had a core radius of 2.1 atomic units for the $s$, $p$ and $d$ channels. In both cases the $d$ channel was used as the local part, and the nonlocal part was converted to the KB form. The self-consistent calculation used a kinetic energy cutoff of 16 Ha for the plane-wave expansion, and a $4\times 4\times 4$ uniform grid for BZ integration (10 $\bf k$} points, once symmetry was taken into account). The lattice constant in the calculations was $5.653\,\,\text{\AA }$.
@@
The calculated effective masses at $\Gamma$ are shown in Table 3. There is an excellent agreement (about six decimal places) between the masses calculated from perturbation theory and those obtained from an interpolation of band energies (with a sixth-order polynomial, and points equally spaced by $10^{-5}$ and centered in $\Gamma$). However, in the interpolation, the bands have to be carefully “disentangled” to follow the correct branch after each crossing. Even with the help of the quality estimate of the fit that flags most mistakes [41], that disentanglement is a very tedious and error-prone job.
A clear sign of the peculiarities in the effective masses is that at $\Gamma$ and in the $[100]$ direction they are identical for the “heavy” and “light” holes. In that direction, the four bands that are degenerate at $\Gamma$ split into two doubly-degenerate bands that are mirror images of each other, and therefore have the same second derivatives at $\Gamma$. From the insets of Fig. 3, it is clear that near $\Gamma$ the bands deviate from parabolas. In Fig. 4, the longitudinal masses along $[100]$ are plotted as a function of distance from the $\Gamma$ point. The reciprocal vector is on a scale $x=\log(1+k/k_{0})$ that is linear near the origin ($k\ll k_{0}$) but logarithmic far from it ($k\gg k_{0}$). The value used in the plot is $k_{0}=10^{-4}\,\,\text{\AA }^{-1}$. The mass of the split-off hole is almost constant throughout the plot range. The heavy-hole and light-hole masses are identical at $\Gamma$, deviating linearly at very small $k$ until they settle at an almost constant value.